OpenStudy (anonymous):

Explain to how to solve for f(3) @ganeshie8 @phi @jdoe0001 @whpalmer4 @KingGeorge

OpenStudy (phi):

f(3) is short for find the formula they give for f(x) replace the x with a 3 in the formula do the arithmetic in the formula and find a number. can you do that ?

OpenStudy (mathmale):

I have helped another person with this problem, and thus know that you have omitted some of the problem statement. Please make sure to share EVERYTHING about each problem when you post a question to OpenStudy;

OpenStudy (mathmale):

Note that you are NOT given f(x); instead, you are given an equation and from it you have to create y our own function f(x).

OpenStudy (anonymous):

@phi for some reason openstudy seemed to be down but can you elabroate on this?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

f(3) means to plug in 3 for x for example if f(x) = x+5 then f(3) = 3+5 = 8

OpenStudy (phi):

you need a formula for f(x) For example, f(x) = 2x + 2 f(3) would mean: replace x with 3: 2*3 + 2 (by the way 2x is short for 2 times x) now simplify. remember order of operations. first do the multiply 2*3=6 . then do the add 6+2 = 8 f(3)= 8

OpenStudy (phi):

Did they give you a formula for f(x) ?

OpenStudy (anonymous):

no i had to write one and cant think of an equation

OpenStudy (phi):

write down the equation for a line. if you really want simple, write y= x

OpenStudy (phi):

or f(x) = x

OpenStudy (mathmale):

@phi: No, they did not provide a formula for f(x). A significant amount of info is missing from this problem statement. What was provided was an EQUATION in x only (no y, no f(x) ). I hope @aleisha96 will double check on the remaining info in this problem statement and provide it to us.

OpenStudy (anonymous):

@mathmale a function was not provided for me i just explained that there is no missing info

OpenStudy (mathmale):

I'm sorry to hear that. I recognize this problem from having worked with another student whose first goal was to find f(3). Looks like there's not much more I can say or do to help if you have no further info.

OpenStudy (anonymous):

@jigglypuff314 can you help me please?

jigglypuff314 (jigglypuff314):

both I and phi have provided examples for your f(3) what more would you like? :)

OpenStudy (anonymous):

i need help forming an equation that will fit into solving for f(3)

jigglypuff314 (jigglypuff314):

alright :) so to do that it would be f(x) = (something) you get to pick what that (something) is :) it can be like x+3 or 2x+5 or even just x

OpenStudy (mathmale):

@aleisha96 : I'd like very much to see you able to finish solving this problem. May I suggest that you plan to ask your teacher, tomorrow, whether there is not more info, and in the meantime move on to solving other problems? Without f(x), you're wandering in the dark (if you'll forgive my putting it that way).

OpenStudy (mathmale):

Or, there could be a ridiculously easy verbal response: "Given the function f(x), evaluate f(3) by replacing every instance of x with 3." That's it. Examples: If f(x) = 5x, f(3) = 5 (3) = 15. If f(x) = sin x, f(3) = sin 3 (assuming that x is an angle and is measured in radians).

OpenStudy (mathmale):

jigglypuff and phi have suggested the same thing: invent f(x) and then substitute 3 for x. Done.

myininaya (myininaya):

Does you problem have anything to do with martians?

OpenStudy (mathmale):

Yes, particularly to do with one Martian named Harry and the other one Richard.